# ADePT: Adaptive Decomposed Prompt Tuning for Parameter-Efficient Fine-tuning

### v2

@@ 1 Introduction
Recently, Pre-trained Large Language Models (PLMs) (Raffel et al., 2020; Touvron et al., 2023) have seen rapid development, with commonly used models now on the scale of hundreds of millions and billions of parameters. Full fine-tuning (FT) of these PLMs requires substantial GPU resources, which is a common challenge faced by both academia and industry. To alleviate this resource-intensive issue, Parameter-Efficient Fine-Tuning (PEFT) (Houlsby et al., 2019; Liu et al., 2022; Hu et al., 2021; Ben Zaken et al., 2022) methods have gained significant attention and have seen breakthrough progress. These PEFT methods tune only a small amount of the internal parameters of a model or extra parameters, allowing PLMs to adapt effectively to target downstream tasks while maintaining performance comparable to FT.
The vanilla Prompt Tuning (PT) (Lester et al., 2021) uses a trainable soft prompt prepended to the input token embeddings (Lester et al., 2021), as shown in Figure 1(a). The few trainable parameters make PT one of the mainstream methods for parameter-efficient fine-tuning. The improvements to PT can be categorized into four paths: the first path involves adding soft prompts to each layer of one PLM (Li & Liang, 2021); the second path involves stabilizing the optimization of soft prompt through a shallow network with a residual connection (Razdaibiedina et al., 2023); the third path involves using soft prompts that had already been trained by other methods for transfer learning (Vu et al., 2022; Asai et al., 2022; Wang et al., 2023); the fourth path uses input token embedding offsets to map the input token embedding into better embedding space (Shi & Lipani, 2024). The first three approaches either require increasing trainable parameters or necessitate additional transfer learning, while the fourth approach requires neither. For the fourth path, Decomposed Prompt Tuning (DePT) (Shi & Lipani, 2024) pioneers the shift in attention from soft prompt to applying input token embeddings offsets to the input token embeddings. As shown in Figure 1(b), DePT learns embedding offsets $\Delta\bm{E}=\bm{A}\bm{B}=[\Delta\bm{e}_{1},\Delta\bm{e}_{2},\cdots,\Delta\bm{e}_{s}]$ by optimizing a pair of low-rank matrices $\bm{A}\in\mathbb{R}^{s\times r_{s}}$ and $\bm{B}\in\mathbb{R}^{r_{s}\times d}$, where $s$ denotes the length of input tokens, $d$ denotes the dimension of input tokens and $r_{s}$ denotes the maximum rank of matrices $\bm{A}$ and $\bm{B}$. Given the input token embeddings $\bm{E}=[\bm{e}_{1},\bm{e}_{2},\cdots,\bm{e}_{s}]\in\mathbb{R}^{s\times d}$, the updated token embeddings $\bm{E}^{\prime}$ are:
@@
DePT also learns a short soft prompt $\bm{P}_{s}\in\mathbb{R}^{l_{s}\times d}$ with length $l_{s}$. The input token embeddings $\bm{E}$, after DePT processing, are formulated as $[\bm{P}_{s},\bm{E}+\bm{A}\bm{B}]$. In this paper, we focus on improving the fourth path by providing better offsets for the input token embeddings. Additionally, our improvement of the fourth path is orthogonal to the first three and can be integrated with their methods.
@@ 2 Related Works
Parameter-Efficient Fine-tuning: PEFT methods adapt PLMs to downstream tasks by optimizing a small number of parameters, significantly reducing computational resource demands. PEFT methods can be divided into four types:(1) Adapter and its variants (Houlsby et al., 2019; He et al., 2022; Rücklé et al., 2021; Ivison & Peters, 2022); (2) Low-rank Adaptation (LoRA) and its variants (Hu et al., 2021; Liu et al., 2022; Kopiczko et al., 2024); (3) Prompt Tuning (PT) and its variants (Lester et al., 2021; Li & Liang, 2021; Vu et al., 2022; Asai et al., 2022; Wang et al., 2023; Ma et al., 2022; Xiao et al., 2023; Razdaibiedina et al., 2023; Shi & Lipani, 2024); (4) other methods (Ben Zaken et al., 2022; Guo et al., 2021). The adapter inserts new modules into the transformer blocks (Houlsby et al., 2019). Hyperformer (He et al., 2022) uses a shared hypernetwork to generate task-conditioned adapters, reducing the trainable parameters in multi-task scenarios. AdapterDrop (Rücklé et al., 2021) removes adapters from lower transformer layers during training and inference, which reduces the computation overhead. Hyperdecoders (Ivison & Peters, 2022) generate input-specific adapters using a shared decoder for multi-task scenarios. LoRA (Houlsby et al., 2019) adopts the matrix product of a pair of low-rank matrices to approximate the updates of corresponding parameters. $(\text{IA})^{3}$ (Liu et al., 2022) rescales internal activations only using learned vectors injected into the attention and feedforward modules. Prompt tuning (PT) (Lester et al., 2021)adapts PLMs to the new downstream tasks by optimizing learnable virtual tokens. LST (Sung et al., 2022) reduces memory requirements by training a small side network, achieving adaptation without back-propagating through the main network. BitFit (Ben Zaken et al., 2022) only tunes the bias of PLMs, which extremely reduces the trainable parameters. Diff pruning learns a sparse “diff” vector to modify a small percentage of pre-trained parameters, enabling efficient adaptation (Guo et al., 2021). Our proposed ADePT is an improved variant of PT, which can also be unified with other PEFT methods.
@@
Prompt tuning and its variants: PT (Lester et al., 2021) enables the adaptation of PLMs to downstream tasks by learning a soft prompt appended in front of model input. Prefix-tuning (Li & Liang, 2021) can be viewed as an extension of Prompt Tuning applied across the entire depth of the model. SPoT (Vu et al., 2022) utilizes trained prompts from source tasks as initialization, but it requires extensive search to find the optimal initialization. ATTEMPT (Asai et al., 2022) adopts an attention module to compose the trained prompts of source tasks. MPT (Wang et al., 2023) learns a transferable prompt by distilling from multiple task-specific prompts. XPrompt (Ma et al., 2022) empirically demonstrates the negative impact of trained prompt tokens and proposes a hierarchical structured pruning for a trained soft prompt, which can be seen as a post-training method. DPT (Xiao et al., 2023) adopts the product of two low-rank matrices to approximate the soft prompt. Residual Prompt Tuning (Razdaibiedina et al., 2023) reparameterizes soft prompt embeddings with a shallow feed-forward network and residual connection, while ADePT uses the feed-forward network to learn input token embedding offsets—a distinct objective. DePT (Shi & Lipani, 2024) first discovers that the input token embedding offsets can enhance the performance of PT and uses short soft prompts to accelerate inference. ADePT produces adaptive input token embedding offsets via shallow token-shared networks, addressing the limitations of DePT mentioned in Section 1. With a comparable number of trainable parameters, ADePT can outperform both vanilla PT and DePT.
@@ 3.2 The Limitations of DePT
Let $\bm{A}\bm{B}=\Delta\bm{E}$, the updated input token embeddings of DePT is formulated as $[\bm{e}_{1}+\Delta\bm{e}_{1},\bm{e}_{2}+\Delta\bm{e}_{2},\cdots,\bm{e}_{s}+\Delta\bm{e}_{s}]$. For a downstream task, the input token embeddings $\bm{E}$ vary. However, the pair of low-rank matrices are fixed after adaptation, meaning that the input token embedding offsets $\Delta\bm{E}$ are position-based. Assume there are input token embeddings $[\bm{a},\bm{b},\bm{c}]$ (here, we omit the right padding), the updated token embeddings are formulated as $[\bm{a}+\Delta\bm{e}_{1},\bm{b}+\Delta\bm{e}_{2},\bm{c}+\Delta\bm{e}_{3}]$. Assume we have a meaningless sequence that does not affect the prediction performance, denoted as $[\bm{t}_{1},\bm{t}_{2}]$. Prepending the meaningless sequence to input token embeddings $[\bm{a},\bm{b},\bm{c}]$, the updated input token embeddings are $[\bm{t}_{1}+\Delta\bm{e}_{1},\bm{t}_{2}+\Delta\bm{e}_{2},\bm{a}+\Delta\bm{e}_{3},\bm{b}+\Delta\bm{e}_{4},\bm{c}+\Delta\bm{e}_{5}]$. We can observe that offsets of $[\bm{a},\bm{b},\bm{c}]$ are different. The offsets of the former is $[\Delta\bm{e}_{1},\Delta\bm{e}_{2},\Delta\bm{e}_{3}]$, while the offsets of the latter is $[\Delta\bm{e}_{3},\Delta\bm{e}_{4},\Delta\bm{e}_{5}]$. Just adding a meaningless sequence $[\bm{t}_{1},\bm{t}_{2}]$ that does not affect the prediction results, the token embedding offsets for $[\bm{a},\bm{b},\bm{c}]$ become different. These position-dependent embedding offsets make the embeddings of each token non-unique in one task, which may be detrimental to the model performance. To simulate this scenario, we design an ideal experiment where we cyclically left shift the column vectors of $\Delta\bm{E}$ and then test the model performance, as shown in Table 1. Let $\Delta\bm{E}^{\prime}$ be the cyclic left shift of $\Delta{\bm{E}}$ by $j$ positions, defined as $\Delta\bm{E}^{\prime}=[\Delta\bm{e}_{1+j},\Delta\bm{e}_{2+j},\cdots,\Delta\bm{e}_{s},\bm{e}_{1},\cdots,\bm{e}_{j}].$ For example, on RTE, the performance of DePT is decreased by 0.7 points after cyclically left-shifting the position-based embedding offsets by 50 positions, and increased by 0.8 points after shifting by 100 positions. On BoolQ, the performance of DePT is worse than the original after cyclically left-shifting the position-based embedding offsets. Table 1 shows that position-based embedding offsets in DePT can cause unstable prediction performance across different positions.
@@ 3.3 Our method: Adaptive Decomposed Prompt Tuning (ADePT)
The limitations of DePT lie in its position-based token embedding offsets. To address this issue, the input token embedding offsets should be tailored for model input, and the corresponding embedding for each token should be unique after being offset. We find that making the input token embedding offsets $\Delta\bm{E}$ a function of the input token embedding $\bm{E}$ can meet such requirements. For an input token embedding $\bm{e}$, we want to get a function $f()$, which can produce the offset of this input token embedding, namely, $f(\bm{e})$. For the input token embeddings $\bm{E}=[\bm{e}_{1},\bm{e}_{2},\cdots,\bm{e}_{s}]$, the updated input token embedding can be formulated as $\bm{E}^{\prime}=\bm{E}+\Delta\bm{E}=[\bm{e}_{1}+f(\bm{e}_{1}),\bm{e}_{2}+f(\bm{e}_{s}),\cdots,\bm{e}_{s}+f(\bm{e}_{s})]$. For example, the input token embeddings $[\bm{a},\bm{b},\bm{c}]$ can be updated as $[\bm{a}+f(\bm{a}),\bm{b}+f(\bm{b}),\bm{c}+f(\bm{c})]$. Prepending the meaningless sequence $[\bm{t}_{1},\bm{t}_{2}]$ to input token embeddings $[\bm{a},\bm{b},\bm{c}]$, the updated input token embeddings are $[\bm{t}_{1}+f(\bm{t}_{1}),\bm{t}_{2}+f(\bm{t}_{2}),\bm{a}+f(\bm{a}),\bm{b}+f(\bm{b}),\bm{c}+f(\bm{c})]$. We can observe that the offsets for $[\bm{a},\bm{b},\bm{c}]$ are the same in this two scenario, i.e., $[f(\bm{a}),f(\bm{b}),f(\bm{c})]$. Therefore, if such a function $f$ exists, we can achieve input token embedding offsets that are tailored for model input, and the embedding for each token is unique within a task.
@@ 3.4 Theoretical analysis
In this section, inspired by Petrov et al. (2024), we provide a theoretical analysis towards ADePT.
@@
$&\bm{o}_{i}^{\text{PT}}=\text{Attention}\left(\bm{e}_{i}\bm{W}_{Q}^{\mathcal{H}},\text{concat}[\bm{P},\bm{E}]\bm{W}_{K}^{\mathcal{H}},\text{concat}[\bm{P},\bm{E}]\bm{W}_{V}^{\mathcal{H}}\right)\\ &=\underbrace{\sum_{k=1}^{l}\bm{A}_{ik}\bm{p}_{k}\bm{W}_{V}^{\mathcal{H}}}_{bias}+\underbrace{\left(1-\sum_{k=1}^{l}\bm{A}_{ik}\right)}_{scale}\bm{o}_{i},\\ \text{with}\quad\bm{A}_{ik}&=\frac{\exp\left(\bm{e}_{i}\bm{W}_{Q}^{\mathcal{H}}\left(\bm{p}_{k}\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)}{\sum_{k=1}^{l}\exp\left(\bm{e}_{i}\bm{W}_{Q}^{\mathcal{H}}\left(\bm{p}^{k}\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)+\sum_{j=1}^{s}\exp\left(\bm{e}_{i}\bm{W}_{Q}^{\mathcal{H}}\left(\bm{e}_{j}\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)},\end{aligned}$ (6)
where $\bm{A}_{ik}$ is the attention score assigned to the prefix vector $\bm{p}_{k}$ for $\bm{e}_{i}$. Thus, in the first transformer layer, PT cannot affect the relative attention patterns across the content and it only scales the attention patterns down while adding a constant bias to the original output $\bm{o}_{i}$ (Petrov et al., 2024).
@@
$&\bm{o}_{i}^{\text{ADePT}}=\text{Attention}\left(\left(\bm{e}_{i}+f\left(\bm{e}_{i}\right)\right)\bm{W}_{Q}^{\mathcal{H}},\text{concat}[\bm{P},\bm{E}+f\left(\bm{E}\right)]\bm{W}_{K}^{\mathcal{H}},\text{concat}[\bm{P},\bm{E}+f\left(\bm{E}\right)]\bm{W}_{V}^{\mathcal{H}}\right)\\ &\quad\quad=\sum_{k=1}^{l}\bm{A}_{ik}\bm{p}_{k}\bm{W}_{V}^{\mathcal{H}}\\ &\quad\quad+(1-\sum_{k=1}^{l}\bm{A}_{ik})\text{Softmax}\left(\left(\left(\bm{e}_{i}+f\left(\bm{e}_{i}\right)\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\left(\bm{E}+f\left(\bm{E}\right)\right)\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)\left(\bm{E}+f\left(\bm{E}\right)\right)\bm{W}_{V}^{\mathcal{H}},\\ &\text{with}\quad\bm{A}_{ik}=\frac{\exp\left(\left(\left(\bm{e}_{i}+f\left(\bm{e}_{i}\right)\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\bm{p}_{k}\bm{W}_{K}^{\mathcal{H}}\right)^{\top}\right)}{B},\\ &\quad\quad B=\sum_{k=1}^{l}\exp\left(\left(\left(\bm{e}_{i}+f\left(\bm{e}_{i}\right)\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\bm{p}_{k}\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)\\ &\quad\quad\quad+\sum_{j=1}^{s}\exp\left(\left(\left(\bm{e}_{i}+f\left(\bm{e}_{i}\right)\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\left(\bm{e}_{j}+f(\bm{e}_{j})\right)\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right).\end{aligned}$ (7)

### v3

@@ 1 Introduction
Recently, Pre-trained Large Language Models (PLMs) (t5; touvron2023llama) have seen rapid development, with commonly used models now on the scale of hundreds of millions and billions of parameters. Full fine-tuning (FT) of these PLMs requires substantial GPU resources, which is a common challenge faced by both academia and industry. To alleviate this resource-intensive issue, Parameter-Efficient Fine-Tuning (PEFT) (houlsby2019parameter; NEURIPS2022_0cde695b; hu2021lora; zaken2021bitfit) methods have gained significant attention and have seen breakthrough progress. These PEFT methods tune only a small amount of the internal parameters of a model or extra parameters, allowing PLMs to adapt effectively to target downstream tasks while maintaining performance comparable to FT.
The vanilla Prompt Tuning (PT) (lester-etal-2021-power) uses a trainable soft prompt prepended to the input token embeddings (lester-etal-2021-power), as shown in Figure 1(a). The few trainable parameters make PT one of the mainstream methods for parameter-efficient fine-tuning. The improvements to PT can be categorized into four paths: the first path involves adding soft prompts to each layer of one PLM (li-liang-2021-prefix); the second path involves stabilizing the optimization of soft prompt through a shallow network with a residual connection (ResidualPrompt); the third path involves using soft prompts that had already been trained by other methods for transfer learning (vu-etal-2022-spot; asai-etal-2022-attempt; MPT); the fourth path uses input token embedding offsets to map the input token embedding into better embedding space (shidept). The first three approaches either require increasing trainable parameters or necessitate additional transfer learning, while the fourth approach requires neither. For the fourth path, Decomposed Prompt Tuning (DePT) (shidept) pioneers the shift in attention from soft prompt to applying input token embeddings offsets to the input token embeddings. As shown in Figure 1(b), DePT learns embedding offsets $\Delta\bm{E}=\bm{A}\bm{B}=[\Delta\bm{e}_{1},\Delta\bm{e}_{2},\cdots,\Delta\bm{e}_{s}]$ by optimizing a pair of low-rank matrices $\bm{A}\in\mathbb{R}^{s\times r_{s}}$ and $\bm{B}\in\mathbb{R}^{r_{s}\times d}$, where $s$ denotes the length of input tokens, $d$ denotes the dimension of input tokens and $r_{s}$ denotes the maximum rank of matrices $\bm{A}$ and $\bm{B}$. Given the input token embeddings $\bm{E}=[\bm{e}_{1},\bm{e}_{2},\cdots,\bm{e}_{s}]\in\mathbb{R}^{s\times d}$, the updated token embeddings $\bm{E}^{\prime}$ are:
@@
DePT also learns a short soft prompt $\bm{P}_{s}\in\mathbb{R}^{l_{s}\times d}$ with length $l_{s}$. The input token embeddings $\bm{E}$, after DePT processing, are formulated as $\bm{P}_{s},\bm{E}+\bm{A}\bm{B}]$. In this paper, we focus on improving the fourth path by providing better offsets for the input token embeddings. Additionally, our improvement of the fourth path is orthogonal to the first three and can be integrated with their methods.
@@ 2 Related Works
Parameter-Efficient Fine-tuning: PEFT methods adapt PLMs to downstream tasks by optimizing a small number of parameters, significantly reducing computational resource demands. PEFT methods can be divided into four types:(1) Adapter and its variants (houlsby2019parameter; he2022hyperprompt; ruckle-etal-2021-adapterdrop; ivison-peters-2022-hyperdecoders); (2) Low-rank Adaptation (LoRA) and its variants (hu2021lora; NEURIPS2022_0cde695b; kopiczko2024vera); (3) Prompt Tuning (PT) and its variants (lester-etal-2021-power; li-liang-2021-prefix; vu-etal-2022-spot; asai-etal-2022-attempt; MPT; ma-etal-2022-xprompt; xiao-etal-2023-decomposed; ResidualPrompt; shidept); (4) other methods (zaken2021bitfit; guo2021diff). The adapter inserts new modules into the transformer blocks (houlsby2019parameter). Hyperformer (he2022hyperprompt) uses a shared hypernetwork to generate task-conditioned adapters, reducing the trainable parameters in multi-task scenarios. AdapterDrop (ruckle-etal-2021-adapterdrop) removes adapters from lower transformer layers during training and inference, which reduces the computation overhead. Hyperdecoders (ivison-peters-2022-hyperdecoders) generate input-specific adapters using a shared decoder for multi-task scenarios. LoRA (houlsby2019parameter) adopts the matrix product of a pair of low-rank matrices to approximate the updates of corresponding parameters. $(\text{IA})^{3}$ (NEURIPS2022_0cde695b) rescales internal activations only using learned vectors injected into the attention and feedforward modules. Prompt tuning (PT) (lester-etal-2021-power)adapts PLMs to the new downstream tasks by optimizing learnable virtual tokens. LST (sunglst) reduces memory requirements by training a small side network, achieving adaptation without back-propagating through the main network. BitFit (zaken2021bitfit) only tunes the bias of PLMs, which extremely reduces the trainable parameters. Diff pruning learns a sparse “diff” vector to modify a small percentage of pre-trained parameters, enabling efficient adaptation (guo2021diff). Our proposed ADePT is an improved variant of PT, which can also be unified with other PEFT methods.
@@
Prompt tuning and its variants: PT (lester-etal-2021-power) enables the adaptation of PLMs to downstream tasks by learning a soft prompt appended in front of model input. Prefix-tuning (li-liang-2021-prefix) can be viewed as an extension of Prompt Tuning applied across the entire depth of the model. SPoT (vu-etal-2022-spot) utilizes trained prompts from source tasks as initialization, but it requires extensive search to find the optimal initialization. ATTEMPT (asai-etal-2022-attempt) adopts an attention module to compose the trained prompts of source tasks. MPT (MPT) learns a transferable prompt by distilling from multiple task-specific prompts. XPrompt (ma-etal-2022-xprompt) empirically demonstrates the negative impact of trained prompt tokens and proposes a hierarchical structured pruning for a trained soft prompt, which can be seen as a post-training method. DPT (xiao-etal-2023-decomposed) adopts the product of two low-rank matrices to approximate the soft prompt. Residual Prompt Tuning (ResidualPrompt) reparameterizes soft prompt embeddings with a shallow feed-forward network and residual connection, while ADePT uses the feed-forward network to learn input token embedding offsets—a distinct objective. DePT (shidept) first discovers that the input token embedding offsets can enhance the performance of PT and uses short soft prompts to accelerate inference. ADePT produces adaptive input token embedding offsets via shallow token-shared networks, addressing the limitations of DePT mentioned in Section 1. With a comparable number of trainable parameters, ADePT can outperform both vanilla PT and DePT.
@@ 3.2 The Limitations of DePT
Let $\bm{A}\bm{B}=\Delta\bm{E}$, the updated input token embeddings of DePT is formulated as $\bm{e}_{1}+\Delta\bm{e}_{1},\bm{e}_{2}+\Delta\bm{e}_{2},\cdots,\bm{e}_{s}+\Delta\bm{e}_{s}]$. For a downstream task, the input token embeddings $\bm{E}$ vary. However, the pair of low-rank matrices are fixed after adaptation, meaning that the input token embedding offsets $\Delta\bm{E}$ are position-based. Assume there are input token embeddings $\bm{a},\bm{b},\bm{c}]$ (here, we omit the right padding), the updated token embeddings are formulated as $\bm{a}+\Delta\bm{e}_{1},\bm{b}+\Delta\bm{e}_{2},\bm{c}+\Delta\bm{e}_{3}]$. Assume we have a meaningless sequence that does not affect the prediction performance, denoted as $\bm{t}_{1},\bm{t}_{2}]$. Prepending the meaningless sequence to input token embeddings $\bm{a},\bm{b},\bm{c}]$, the updated input token embeddings are $\bm{t}_{1}+\Delta\bm{e}_{1},\bm{t}_{2}+\Delta\bm{e}_{2},\bm{a}+\Delta\bm{e}_{3},\bm{b}+\Delta\bm{e}_{4},\bm{c}+\Delta\bm{e}_{5}]$. We can observe that offsets of $\bm{a},\bm{b},\bm{c}]$ are different. The offsets of the former is $\Delta\bm{e}_{1},\Delta\bm{e}_{2},\Delta\bm{e}_{3}]$, while the offsets of the latter is $\Delta\bm{e}_{3},\Delta\bm{e}_{4},\Delta\bm{e}_{5}]$. Just adding a meaningless sequence $\bm{t}_{1},\bm{t}_{2}]$ that does not affect the prediction results, the token embedding offsets for $\bm{a},\bm{b},\bm{c}]$ become different. These position-dependent embedding offsets make the embeddings of each token non-unique in one task, which may be detrimental to the model performance. To simulate this scenario, we design an ideal experiment where we cyclically left shift the column vectors of $\Delta\bm{E}$ and then test the model performance, as shown in Table 1. Let $\Delta\bm{E}^{\prime}$ be the cyclic left shift of $\Delta{\bm{E}}$ by $j$ positions, defined as $\Delta\bm{E}^{\prime}=[\Delta\bm{e}_{1+j},\Delta\bm{e}_{2+j},\cdots,\Delta\bm{e}_{s},\bm{e}_{1},\cdots,\bm{e}_{j}].$ For example, on RTE, the performance of DePT is decreased by 0.7 points after cyclically left-shifting the position-based embedding offsets by 50 positions, and increased by 0.8 points after shifting by 100 positions. On BoolQ, the performance of DePT is worse than the original after cyclically left-shifting the position-based embedding offsets. Table 1 shows that position-based embedding offsets in DePT can cause unstable prediction performance across different positions.
@@ 3.3 Our method: Adaptive Decomposed Prompt Tuning (ADePT)
The limitations of DePT lie in its position-based token embedding offsets. To address this issue, the input token embedding offsets should be tailored for model input, and the corresponding embedding for each token should be unique after being offset. We find that making the input token embedding offsets $\Delta\bm{E}$ a function of the input token embedding $\bm{E}$ can meet such requirements. For an input token embedding $\bm{e}$, we want to get a function $f()$, which can produce the offset of this input token embedding, namely, $f(\bm{e})$. For the input token embeddings $\bm{E}=[\bm{e}_{1},\bm{e}_{2},\cdots,\bm{e}_{s}]$, the updated input token embedding can be formulated as $\bm{E}^{\prime}=\bm{E}+\Delta\bm{E}=[\bm{e}_{1}+f(\bm{e}_{1}),\bm{e}_{2}+f(\bm{e}_{s}),\cdots,\bm{e}_{s}+f(\bm{e}_{s})]$. For example, the input token embeddings $\bm{a},\bm{b},\bm{c}]$ can be updated as $\bm{a}+f(\bm{a}),\bm{b}+f(\bm{b}),\bm{c}+f(\bm{c})]$. Prepending the meaningless sequence $\bm{t}_{1},\bm{t}_{2}]$ to input token embeddings $\bm{a},\bm{b},\bm{c}]$, the updated input token embeddings are $\bm{t}_{1}+f(\bm{t}_{1}),\bm{t}_{2}+f(\bm{t}_{2}),\bm{a}+f(\bm{a}),\bm{b}+f(\bm{b}),\bm{c}+f(\bm{c})]$. We can observe that the offsets for $\bm{a},\bm{b},\bm{c}]$ are the same in this two scenario, i.e., $f(\bm{a}),f(\bm{b}),f(\bm{c})]$. Therefore, if such a function $f$ exists, we can achieve input token embedding offsets that are tailored for model input, and the embedding for each token is unique within a task.
@@ 3.4 Theoretical analysis
In this section, inspired by promptheory, we provide a theoretical analysis towards ADePT.
@@
$\begin{aligned} &\bm{o}_{i}^{\text{PT}}=\text{Attention}\left(\bm{e}_{i}\bm{W}_{Q}^{\mathcal{H}},\text{concat}[\bm{P},\bm{E}]\bm{W}_{K}^{\mathcal{H}},\text{concat}[\bm{P},\bm{E}]\bm{W}_{V}^{\mathcal{H}}\right)\\ &=\underbrace{\sum_{k=1}^{l}\bm{A}_{ik}\bm{p}_{k}\bm{W}_{V}^{\mathcal{H}}}_{bias}+\underbrace{\left(1-\sum_{k=1}^{l}\bm{A}_{ik}\right)}_{scale}\bm{o}_{i},\\ \text{with}\quad\bm{A}_{ik}&=\frac{\exp\left(\bm{e}_{i}\bm{W}_{Q}^{\mathcal{H}}\left(\bm{p}_{k}\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)}{\sum_{k=1}^{l}\exp\left(\bm{e}_{i}\bm{W}_{Q}^{\mathcal{H}}\left(\bm{p}^{k}\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)+\sum_{j=1}^{s}\exp\left(\bm{e}_{i}\bm{W}_{Q}^{\mathcal{H}}\left(\bm{e}_{j}\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)},\end{aligned}$ (6)
where $\bm{A}_{ik}$ is the attention score assigned to the prefix vector $\bm{p}_{k}$ for $\bm{e}_{i}$. Thus, in the first transformer layer, PT cannot affect the relative attention patterns across the content and it only scales the attention patterns down while adding a constant bias to the original output $\bm{o}_{i}$ (promptheory).
@@
$\begin{aligned} &\bm{o}_{i}^{\text{ADePT}}=\text{Attention}\left(\left(\bm{e}_{i}+f\left(\bm{e}_{i}\right)\right)\bm{W}_{Q}^{\mathcal{H}},\text{concat}[\bm{P},\bm{E}+f\left(\bm{E}\right)]\bm{W}_{K}^{\mathcal{H}},\text{concat}[\bm{P},\bm{E}+f\left(\bm{E}\right)]\bm{W}_{V}^{\mathcal{H}}\right)\\ &\quad\quad=\sum_{k=1}^{l}\bm{A}_{ik}\bm{p}_{k}\bm{W}_{V}^{\mathcal{H}}\\ &\quad\quad+(1-\sum_{k=1}^{l}\bm{A}_{ik})\text{Softmax}\left(\left(\left(\bm{e}_{i}+f\left(\bm{e}_{i}\right)\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\left(\bm{E}+f\left(\bm{E}\right)\right)\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)\left(\bm{E}+f\left(\bm{E}\right)\right)\bm{W}_{V}^{\mathcal{H}},\\ &\text{with}\quad\bm{A}_{ik}=\frac{\exp\left(\left(\left(\bm{e}_{i}+f\left(\bm{e}_{i}\right)\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\bm{p}_{k}\bm{W}_{K}^{\mathcal{H}}\right)^{\top}\right)}{B},\\ &\quad\quad B=\sum_{k=1}^{l}\exp\left(\left(\left(\bm{e}_{i}+f\left(\bm{e}_{i}\right)\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\bm{p}_{k}\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)\\ &\quad\quad\quad+\sum_{j=1}^{s}\exp\left(\left(\left(\bm{e}_{i}+f\left(\bm{e}_{i}\right)\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\left(\bm{e}_{j}+f(\bm{e}_{j})\right)\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right).\end{aligned}$ (7)